OpenStudy (cutiecomittee123):

Pre calculus help! Find the sum of the first six terms of the geometric sequence for which a2=0.7 and a3=0.49

OpenStudy (joannablackwelder):

Do you have the formula for the sum of a geometric sequence?

OpenStudy (cutiecomittee123):

I know the equation you use is Sn=a1(1-r^n)/1-r

OpenStudy (joannablackwelder):

Yep, good. Se we need to identify a1, r, and n. Any ideas on a1?

OpenStudy (cutiecomittee123):

I also know that we have to find the difference between 0.7 and 0.49

OpenStudy (cutiecomittee123):

which will get us to what r is equal to

OpenStudy (cutiecomittee123):

but I am hvai

OpenStudy (joannablackwelder):

Well, r is the common ratio since this is a geometric sequence.

OpenStudy (cutiecomittee123):

having trouble with those calculations because I dont know if its 1/7 or 0.7

OpenStudy (joannablackwelder):

The common ratio can be found using \[\frac{ a _{n+1} }{ a _{n} }\]

OpenStudy (cutiecomittee123):

so 0.49+1/0.7

OpenStudy (joannablackwelder):

No, the +1 is in the subscript. Meaning take the ratio of a term divided by the term before it.

OpenStudy (joannablackwelder):

In this case, .49/.7

OpenStudy (cutiecomittee123):

Oh I see, when I calculate that I get 1/7

OpenStudy (cutiecomittee123):

or .7

OpenStudy (joannablackwelder):

1/7 isn't .7

OpenStudy (joannablackwelder):

7/10 is .7

OpenStudy (cutiecomittee123):

oh then .7

OpenStudy (joannablackwelder):

Right, good. So that is r.

OpenStudy (joannablackwelder):

What do you think a1 is?

OpenStudy (cutiecomittee123):

I think we use a2 which is 0.7 and then divide that by the common ration r=.7 When I did that I got 1 because .7/.7 =1

OpenStudy (joannablackwelder):

Yep, perfect. :-)

OpenStudy (joannablackwelder):

And how about n?

OpenStudy (cutiecomittee123):

Okay and then the equation for the sum is S6=1(1-.7^6)/

OpenStudy (cutiecomittee123):

.7

OpenStudy (cutiecomittee123):

n=6 because thats the numbers we have in the sum

OpenStudy (cutiecomittee123):

well no, the numbers we are using to find the sum

OpenStudy (joannablackwelder):

Right, good. But in my equation, the denominator is 1-.7

OpenStudy (cutiecomittee123):

Okay yeah thats how I plugged it in and I came up with S6=.182351

OpenStudy (joannablackwelder):

Hm, that's not what I get.

OpenStudy (joannablackwelder):